OpenStudy (anonymous):

Match each number in standard notation with its value in exponential form. ]0.01 100 0.0001 1,000 (A.) 10^-4 (B.) 10^3 (C.) 10^-2 (D.) 10^2

OpenStudy (anonymous):

how is this tough? :/ 10^-4 is 0.0001 10^3 is 1000 10^-2 is 0.01 10^2 is 100

OpenStudy (anonymous):

Look at these examples. The first says 10 to the second power, or 10 squared, is 10 times 10 or 100. Then 10 cubed is 10 times 10 times 10 or 1,000. Next, 10 to the fourth is 10 times 10 times 10 times 10 or 10,000. Also, 10 to the first power is simply 10. Less obviously, 10 to the zero power equals one. Note that 10 to any positive power is just one followed by that number of zeros. 10^2 = 10 x 10 = 100 10^3 = 10 x 10 x 10 = 1,000 10^4 = 10 x 10 x 10 x 10 = 10,000 also 10^1 = 10 and 10^0 = 1

OpenStudy (anonymous):

i dnt remeber this stuff im in 12th grade and im helping my lil sis. Its been a while

OpenStudy (anonymous):

i am in 12th grade and i still remember this "Stuff"

OpenStudy (anonymous):

about to join a college too

OpenStudy (anonymous):

i dont have many brain cells left lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Maybe this will refresh you memory: \(10^5=100000\) \(10^4=10000\) \(10^3=1000\) \(10^2=100\) \(10^1=10\) \(10^0=1\) \(10^{-1}=0.1\) \(10^{-2}=0.01\) \(10^{-3}=0.001\) \(10^{-4}=0.0001\) \(10^{-5}=0.00001\) :)

OpenStudy (anonymous):

Thank you both it did help :) @KAOS|abdullah and @salehgeek

OpenStudy (anonymous):

no prob mate ;)

OpenStudy (anonymous):

Happy to help!

OpenStudy (anonymous):

For more insight see: http://www.onlinemathlearning.com/scientific-notation.html